OpenStudy (kevin):

These questions about number pattern is given to semester 1 grade 7 in my country and used as a standard curriculum. Would you try? 1. Specify p = 1/3 + 1/9 + 1/27 + 1/81 + .... 2. Specify y = x + 13 + x + 23 + x + 33 + ... + x + 1003 3. 23a23b is divisible by 8 and 9. Specify a + b 4. If 0,201020102010 .... = x/y. Then the smallest number of x + y is 5.If this day is Tuesday, then 5^2000 day again is 6. Sandy celebrate her 25th anniversary on Friday. If during her lifetime, she spent 6 leap years, then Sandy was born on what day?

OpenStudy (will.h):

1- it goes by 1/3x we multiply the denominator by 3 each time 2- adding 10 to the denominator each time 6 - a leap year occurs every 4 years so she was born she witnessed 6 leap years so there has been 6 day added in her entire life-time to her birthday so Friday - 6 days = Saturday 5 - i have no clue how to do this but i would say: 5^2000 5 represent Tuesday and 2000 represent the number of days in those 2000 days there would be 166.67 years in those years there would be 41.67 leap years so we have 166.67 - 41.67 = 125 original years and 41.67 leaps years lol so tricky i don't even know why i calculated those

jhonyy9 (jhonyy9):

1. let 3 being an x ,than your equation will be 1 1 1 1 p = --- + ----- + ----- + ----- + ...... x x^2 x^3 x^4 or p = x^(-1) +x^(-2) +x^(-3) +x^(-4) + ... 2. y = x+13 +x+23 +x+33+... +x+1003 let x+13 note an p so than your equation will be y = p +p+10 + p +20 + .... + p+990 3. 23a23b divisible by 8 and 9 specify a+b 23^2(a*b)/(8*9) 23^2(8*9)/(8*9) so a=8 and b=9 and a+b = 17 4. will the second have not denominator there are sum of these x+13 and x+23

OpenStudy (will.h):

i actually quit Ciao

jhonyy9 (jhonyy9):

@Will.H these above wrote is to you - that the second have not denominators

jhonyy9 (jhonyy9):

ciao to you to - are you Italian ? just bc. they say CIAO - or Saluti or Arrivederci

OpenStudy (will.h):

@jhonyy9 half Italian lol good knowledge tho

jhonyy9 (jhonyy9):

cosi posso tio parlo italiano ma solo pocco perque i parlo piccolo ,ho laborato 7 anni a una firma taliano qui in hungaria

jhonyy9 (jhonyy9):

ma no lo so te capisi o no mio paroli - sorry

OpenStudy (will.h):

Oh my god too close well we both Europeans lol but i was born in the states so i have limited ability in speaking Italian is Hungarian the same! i think they share a few

jhonyy9 (jhonyy9):

not is problem - hope you understand me with my english too - so this is really true that your parents not learned you nothing from italian lenguage - or your grand parents not told to you in italian - or you dont visit never Italy ?

OpenStudy (will.h):

Been to Italy Rome several times. yeah my grandparents immigrated to USA and my parents did not care about Italian lol but i was curious so i catche up a few words/ expressions and so

OpenStudy (will.h):

I think Kevin will hate us bc we keep spamming the question lol

jhonyy9 (jhonyy9):

hope so much that Kevin will can understanding that my subbmitt button again is blocked from where my past profile of champion was deleted - and again not was desblocked ,activated till today - Kevin sorry for this all - but hope so much that you remember about my blocked submitt button case -

OpenStudy (will.h):

Yup sorry Kevin for spamming lol and jhonny you had another account?

jhonyy9 (jhonyy9):

yes past 4 or 5 years and in this year in month of may was deleted same with other to much profiles - hope you heard about these or you can reading about these on openstudyfeedback posted questions too -

OpenStudy (will.h):

Oh sorry to hear. Your current account looks good tho high level and smart score

OpenStudy (kevin):

Hmm... sorry, but both of you have wrong answers. xD Also for number 1 and 2, you must solve for p and y.

OpenStudy (jiteshmeghwal9):

The first one is a G.P.

OpenStudy (kevin):

what is that?

OpenStudy (jiteshmeghwal9):

The common ratio in the first question is 1/3. Sum= \[\frac{a}{1-r}\]

OpenStudy (jiteshmeghwal9):

G.P.=Geometric progression

OpenStudy (kevin):

Owh.. you mean geometric progression. But, in 7th grade... the teacher still not teach them about geometric yet. You should solve it using number pattern

OpenStudy (kevin):

They don't know about arithmetic and geometric...

OpenStudy (kevin):

Well, forget that xD You can solve it using geometric too actually. How about number 3-6?

OpenStudy (jiteshmeghwal9):

What are the divisibility test for 8 and 9?

OpenStudy (kevin):

yup.. it's for number 3

OpenStudy (jiteshmeghwal9):

Divisibility test for 8= a given number is divisible by 8 if the number formed by its last three digits is divisible by 8

OpenStudy (jiteshmeghwal9):

A given number is divisible by 9 if sum of its digits is divisible by 9

OpenStudy (kevin):

Using 23b is divisible by 8, then you'll find b = 2 Using (2 + 3 + a + 2 + 3 + 2) is divisible by 9, -> (12 + a) is divisible by 9, then you'll find a = 6

OpenStudy (will.h):

jit mate a week ago you were at level 70 lol moving fast

OpenStudy (kevin):

Yeah... that's it jitesh

OpenStudy (kevin):

how about number 4?

OpenStudy (kevin):

He was 79 before Will, I saw it xD

OpenStudy (kevin):

I mean 89

OpenStudy (will.h):

fair enough

OpenStudy (jiteshmeghwal9):

Yes will u r also quicker

OpenStudy (will.h):

Not really lol

OpenStudy (will.h):

sorry for spamming again kevin u 2 keep up

OpenStudy (kevin):

You too will...

OpenStudy (kevin):

Thx

OpenStudy (jiteshmeghwal9):

Lol will Let x=0.2010...... 10000x=2010.2010..... 10000x=2010+x 9999x=2010

OpenStudy (kevin):

For number 1 : p = 1/3 + 1/9 +1/27 + 1/81+ .... p = 1/3 (1 + 1/3 + 1/9 + 1/27 ....) p = 1/3 (1 + p) p = 1/3 +1/3p p - 1/3p = 1/3 2/3p = 1/3 p = 3/6 = 1/3 It's same for number 2 Maybe you'll get an idea for number 4-6 after seeing this...

OpenStudy (kevin):

Yeah.. that's true... lol You are really awesome jitesh xD Last question, how about question 5? Never mind question 6, it's same with question 5